OpenStudy (anonymous):

a ball starts from rest and rolls down a hill with uniform acceleration, traveling 150m in the interval t=3 to t=7, how far did the ball travel in the first 3 seconds?

OpenStudy (anonymous):

whr r u stuck

OpenStudy (anonymous):

let x be the distance traveled in the first 3 sec 1: x+150 = 1/2*a*7^2 (total distance starting at t= 0 V=0) 2: 150 = Vi*4 +1/2*a*4^2 (Vi here means the velocity at t=3) 3: Vi = a*3 (velocity after 3 seconds starting from rest) use eqn 3 to sub.s into 2 to find 'a' use 'a' to sub.s into 1 to find 'x'